Given 3x^2+(k+2)x=-3, determine the value(s) of k so that the equation has real and equal roots.
Answer by stanbon(75887)

Given 3x^2+(k+2)x=-3, determine the value(s) of k so that the equation has real and equal roots.
----------------
The discriminant must be zero.
b^2-4ac = (k+2)^2-4*3*3 = 0
(k+2)^2 = 36
k+2 = 6 or k+2 = -6
k = 4 or k = -8
==================
Cheers,
Stan H.